# GeoSurDepth: Spatial Geometry-Consistent Self-Supervised Depth Estimation for Surround-View Cameras

### v1

@@ IV-A Implementation Details
Training. Our networks were implemented in PyTorch [18] and trained on four NVIDIA RTX 4090 GPUs. MonoViT-Small [19], adapted to our surround-view setting, is employed as the depth network. ResNet-18 [20] was adopted as pose network following VFDepth [6]. During training, images from the previous and subsequent frames ($t^{\prime}\in{t-1,t+1}$) were used as temporal context. We trained the models using the Adam optimizer [21] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, a learning rate of $1\times 10^{-4}$, and 30/20 training epochs for DDAD/nuScenes dataset. A batch size of 1, consisting of images from six cameras, was used per GPU. Focal normalization [22] and the intensity alignment strategy proposed in VFDepth [6] were applied during training. The weighting coefficients of the loss functions were set as $\lambda_{\text{T}}=1$, $\lambda_{\text{S}}=0.03$, $\lambda_{\text{ST}}=0.1$, $\lambda_{\text{MVRC}}=0.2$, $\omega_{p}=1$, $\omega_{s}=0.001$, $\omega_{\text{SDC}}=0.001$, $\omega_{\text{SNC}}=0.01$, $\omega_{\text{DSC}}=1$, $\kappa_{\text{SRC}}=0.1$, $\kappa_{\text{SNC}}=0.1$, $\mu=0.1$. ViT-Base variant of DepthAnything V2 and ViT-B/32 variant of CLIP model were used as foundation models. Self-occlusion masks and reprojection masks were applied to exclude invalid pixels from loss computation.
@@ IV-B Experiment Results
Figure 4: Visualization of depth estimation results on the DDAD (above) and nuScenes (below) datasets. White boxes indicate erroneous estimations. As observed, DepthAnything may also output inaccurate estimations in certain areas. We thus use it indirectly as pseudo priors.
We compare our proposed method with other state-of-the-art approaches. Quantitative evaluations for both metric and scale-ambiguous depth estimation on DDAD and nuScenes datasets are reported in Table I. Our method achieves substantially better or competitive performance under both scale-aware and ambiguous evaluation protocols compared with existing baselines and Depth Anything V2 on both datasets, even when using ResNet34 as depth network. Qualitative visualizations of surround-view depth estimation results on both datasets are presented in Fig.4. As shown, our method produces smooth, edge- and object-aware depth maps, significantly outperforming methods such as CVCDepth. Visualization of more examples as well as point cloud reconstruction effect on both datasets can be found in Supplementary material provided.
TABLE I: Depth evaluation results on DDAD [16] and nuScenes [17] datasets (* indicates reproduced results by VFDepth [6]. -S1 indicates Stage 1 results of GaussianOcc [24]. -M and -S indicate scale-aware and scale-ambiguous variant of SurroundDepth [5], respectively. $\dagger$ indicates ResNet-34+horizontal flip post-process evaluation of CVCDepth [13]. - implies results not reported. Best result in bold, second best underlined).
@@ IV-C Ablation Studies
Adaptive joint motion learning. Table II demonstrates improved depth estimation accuracy for both individual camera views and the surround-view setting. We attribute this improvement to the learnable weighting of features extracted by the pose encoder, which enables the network to adaptively emphasize informative views and thereby achieve more effective structure-from-motion learning.
TABLE II: Ablation study on different motion learning strategies on DDAD dataset (F, B, L, R implies front, back, left and right).
@@
Spatial dense depth-based view synthesis. In this work, we leverage reconstructed spatial dense depth to further perform view synthesis and enforce surface normal consistency. As shown in Table III, removing either $\tilde{\mathcal{L}}_{\text{SRC}}$ or $\tilde{\mathcal{L}}_{\text{SNC}}$ from the overall loss function leads to noticeable performance degradation. Fig.5 presents a qualitative example of view synthesis using both the estimated depth and the reconstructed spatial dense depth through spatial warping. The zoomed-in regions illustrate the complementary effect of these two depth sources on view synthesis and photometric penalization for 3D lifting.
TABLE III: Evaluation of the use of reconstructed spatial dense depth on metric depth estimation of DDAD dataset.
@@
Geometry guidance and enhancement via foundation models. Results reported in Table IV demonstrate that enforcing 3D geometric accuracy through surface normal consistency, together with regularizing edge-aware depth estimation using 2D disparity gradients, leads to improved model performance. In contrast, removing CLIP from the depth network results in a slight drop in model performance. Fig.6 provides qualitative comparisons of these variants, illustrating less smooth depth transitions in low-texture regions, increased edge blurring at object boundaries when geometric guidance is absent, and reduced semantic detail when geometric feature representation enhancement is removed.
TABLE IV: Ablation study on geometry guidance by DepthAnything and feature representation enhancement by CLIP model.
@@
Meanwhile, we also conduct experiments using different encoder variants of DepthAnything V2. As shown in Table V, the different encoders exhibit largely comparable performance, with ViT-B achieving slightly better results. We attribute this to the relatively low input image resolution used by the network, under which fine-grained details may be compressed, limiting the advantage of more powerful encoders.
TABLE V: Comparison of metric depth estimation results with different encoder of DepthAnything V2 on DDAD dataset.
[…]
ViT-S 0.177 2.782 11.639 0.284 0.759 0.909 0.955
ViT-B 0.176 2.738 11.520 0.280 0.763 0.912 0.957
ViT-L 0.175 2.682 11.895 0.284 0.754 0.906 0.955
@@ V Conclusion
In this work, we presented GeoSurDepth, a framework for self-supervised surround-view depth estimation that exploits geometry consistency as primary cue. By integrating foundation models as pseudo-geometry priors and for feature enhancement, enforcing 3D surface normal consistency, and regularizing object- and texture-level depth, GeoSurDepth achieves accurate and edge-aware depth estimations. A novel view synthesis pipeline provides additional photometric supervision through 2D-3D lifting and multi-contextual reconstruction, while an adaptive joint motion learning strategy enables the network to emphasize informative camera views for improved motion reasoning. Extensive experiments on DDAD and nuScenes demonstrate that GeoSurDepth achieves state-of-the-art performance, highlighting the importance of exploiting geometry coherence and consistency for robust multi-view depth estimation.
VI Supplementary material
VI-A Method
VI-A1 Pseudo depth of DepthAnything
We denote the inverse depth representation directly output by DepthAnything as $\mathbf{S}^{\text{DA}}$. Due to the inherent scale and shift ambiguity of monocular depth estimation, this representation cannot be directly interpreted as a metric quantity. Instead, it can be formulated as an affine transformation of the true inverse depth,
$\mathbf{S}^{\text{DA}}=\alpha\frac{1}{\mathbf{D}^{\text{true}}}+\beta,$ (28)
where $\mathbf{D}^{\text{true}}$ denotes the true depth in the physical world, and $\alpha$ and $\beta$ are unknown scale and offset parameters.
To remove this affine ambiguity, we apply min-max normalization to $\mathbf{S}^{\text{DA}}$,
$\bar{\mathbf{S}}^{\text{DA}}=\frac{\mathbf{S}^{\text{DA}}-\mathbf{S}^{\text{DA}}_{\text{min}}}{\mathbf{S}^{\text{DA}}_{\text{max}}-\mathbf{S}^{\text{DA}}_{\text{min}}}.$ (29)
Substituting (28) into (29) gives
$\bar{\mathbf{S}}^{\text{DA}}=\frac{\frac{\alpha}{\mathbf{D}^{\text{true}}}+\beta-\left(\frac{\alpha}{\mathbf{D}^{\text{true}}_{\text{max}}}+\beta\right)}{\left(\frac{\alpha}{\mathbf{D}^{\text{true}}_{\text{min}}}+\beta\right)-\left(\frac{\alpha}{\mathbf{D}^{\text{true}}_{\text{max}}}+\beta\right)}=\frac{\frac{1}{\mathbf{D}^{\text{true}}}-\frac{1}{\mathbf{D}^{\text{true}}_{\text{max}}}}{\frac{1}{\mathbf{D}^{\text{true}}_{\text{min}}}-\frac{1}{\mathbf{D}^{\text{true}}_{\text{max}}}},$ (30)
which shows that the normalized representation $\bar{\mathbf{S}}^{\text{DA}}$ is invariant to the unknown affine parameters $\alpha$ and $\beta$, and depends solely on the relative inverse-depth distribution.
We further interpret $\bar{\mathbf{S}}^{\text{DA}}$ as a normalized disparity-like representation and map it to a predefined target depth range $\mathbf{D}^{\text{tgt}}_{\text{min}},\mathbf{D}^{\text{tgt}}_{\text{max}}]$. Specifically, we define
$\text{disp}_{\text{min}}=\frac{1}{\mathbf{D}^{\text{tgt}}_{\text{max}}},~\text{disp}_{\text{max}}=\frac{1}{\mathbf{D}^{\text{tgt}}_{\text{min}}},$ (31)
and recover the estimated depth as,
$\mathbf{D}^{\text{DA}}$ $=\frac{1}{\text{disp}_{\text{min}}+(\text{disp}_{\text{max}}-\text{disp}_{\text{min}})\cdot\bar{\mathbf{S}}^{\text{DA}}}$ (32)
$=\frac{1}{\frac{1}{\mathbf{D}^{\text{tgt}}_{\text{max}}}+\left(\frac{1}{\mathbf{D}^{\text{tgt}}_{\text{min}}}-\frac{1}{\mathbf{D}^{\text{tgt}}_{\text{max}}}\right)\cdot\bar{\mathbf{S}}^{\text{DA}}}$
$=\frac{1}{\frac{1}{\mathbf{D}^{\text{tgt}}_{\text{max}}}+\left(\frac{1}{\mathbf{D}^{\text{tgt}}_{\text{min}}}-\frac{1}{\mathbf{D}^{\text{tgt}}_{\text{max}}}\right)\cdot\frac{1/\mathbf{D}^{\text{true}}-1/\mathbf{D}^{\text{true}}_{\text{max}}}{1/\mathbf{D}^{\text{true}}_{\text{min}}-1/\mathbf{D}^{\text{true}}_{\text{max}}}}.$
In practice, both $\mathbf{D}^{\text{true}}_{\text{max}}$ and $\mathbf{D}^{\text{tgt}}_{\text{max}}$ are typically large, such that their reciprocals can be approximated as zero. Under this approximation, the above expression simplifies to
$\mathbf{D}^{\text{DA}}\approx\frac{1/\mathbf{D}^{\text{true}}_{\text{min}}-1/\mathbf{D}^{\text{true}}_{\text{max}}}{1/\mathbf{D}^{\text{tgt}}_{\text{min}}-1/\mathbf{D}^{\text{tgt}}_{\text{max}}}\cdot\mathbf{D}^{\text{true}}=\frac{1}{\gamma}\mathbf{D}^{\text{true}},$ (33)
where $\gamma$ is a scale factor that characterizes the proportional relationship between the pseudo depth inferred from DepthAnything and the true depth in the physical world. Based on this property, we further compute the surface normal vector map from the pseudo depth by DepthAnything, and reconstruct spatial dense depth through cross-view geometry.
VI-A2 Spatial dense depth reconstruction methods
In this work, we adopt the modified spatial backward warping strategy proposed in CVCDepth [13] to reconstruct spatial dense depth from adjacent views. We further summarize representative depth reprojection and reconstruction strategies employed in prior studies.
(1) Forward warping (FW) [26]. FW lifts each pixel from source view into 3D space using the estimated depth, and transforms it into target camera coordinate system via extrinsics. The transformed 3D point is then projected onto the target image plane, where its depth value is assigned to corresponding pixel to form warped depth map. FW is geometrically correct. However, it does not define a one-to-one mapping and also results in holes in depth map due to discretization. FW can be formulated as,
$\mathbf{P}_{j}=\hat{\mathbf{D}}_{j}(\mathbf{p}_{j})\mathbf{K}_{j}^{-1}\mathbf{p}_{j},$ (34)
$\tilde{\mathbf{P}}_{j}=\mathbf{E}_{i}\mathbf{E}_{j}^{-1}\mathbf{P}_{j},~\tilde{\mathbf{D}}_{j}(\mathcal{Q}(\mathbf{K}_{i}\tilde{\mathbf{P}}_{j}))=(\tilde{\mathbf{P}}_{j})_{z},$ (35)
where $\tilde{\mathbf{P}}_{j}\in\mathbb{R}^{3}$ denotes a 3D point reprojected from camera $j$ to the coordinate frame of camera $i$. $(\cdot)_{z}$ implies $z$ value of a point cloud. $\mathcal{Q}(\cdot)$ denotes discretization operator that maps continuous coordinates to pixel indices.
(2) Backward warping (BW). BW follows the same synthesis procedure as bilinear-sampling-based view synthesis with only source frame of a depth map. It does not provide correct depth reference as same objects have different depth value in different viewpoints. BW can be formulated as follows.
$\mathbf{p}_{ij}=\mathbf{\Pi}_{ij}\mathbf{p}_{i},~\tilde{\mathbf{D}}_{j}(\mathbf{p})=\hat{\mathbf{D}}_{j}\left<\mathbf{p}_{ij}\right$> (36)
(3) Modified backward warping (MBW). To deal with issue of BW, CVCDepth [13] implements bilinear sampling on source depth map transformed to target view to facilitate geometry correctness. MBW can be formulated as follows.
$\tilde{\mathbf{D}}_{j}(\mathbf{p})=(\tilde{\mathbf{P}}_{j})_{z}\left<\mathbf{p}_{i\rightarrow j}\right$> (37)
(4) Modified forward+backward warping (MFBW). MFBW, proposed and used in MonoDiffusion [27] and [28], constructs depth filtering masks for knowledge distillation from a teacher network. It first obtains a depth map via BW and uses it to lift 2D pixels into 3D space, followed by a 3D-2D projection as in FW to deal with discretization issue. However, the reprojection relies on bilinearly sampled depth values on the source view, which may correspond to interpolated 3D points that do not strictly exist in the physical scene. MFBW can be formulated as follows,
$\mathbf{P}_{j}^{\text{BW}}=\hat{\mathbf{D}}_{j}^{\text{BW}}(\mathbf{p}_{j})\mathbf{K}_{j}^{-1}\mathbf{p}_{j},$ (38)
$\tilde{\mathbf{P}}_{j}^{\text{BW}}=\mathbf{E}_{i}\mathbf{E}_{j}^{-1}\mathbf{P}_{j},~\tilde{\mathbf{D}}_{j}(\mathbf{p})=(\tilde{\mathbf{P}}_{j}^{\text{BW}})_{z}.$ (39)
VI-A3 Proof of scale- and shift-invariance in surface normal vector map computation with reconstructed spatial dense depth of DepthAnything
In this work, both $\mathcal{L}_{\text{SNC}}$ and $\tilde{\mathcal{L}}_{\text{SNC}}$ are adopted as components of the overall loss function. In the submitted main manuscript, we present a derivation demonstrating the scale invariance of surface normal vector computation when using the scale-ambiguous depth produced by DepthAnything. Here, we provide a detailed derivation showing that the surface normal vector map computed from reconstructed spatial dense depth of DepthAnything is also invariant to scale and shift.
Specifically, we perform modified spatial backward warping on the pseudo scale-ambiguous depth by DA as follows.
$\mathbf{P}_{j}=D_{j}^{\text{true}}\mathbf{K}^{-1}_{j}\mathbf{p}_{j}\approx\gamma D_{j}^{\text{DA}}\mathbf{K}_{j}^{-1}\mathbf{p}_{j},$ (40)
$\tilde{D}_{j}^{\text{DA}}=\left<(\tilde{\mathbf{P}}_{j})_{z}\right>_{\mathbf{p}_{ij}}=\left<(\mathbf{R}_{ij}\mathbf{P}_{j}+\mathbf{t}_{ij})_{z}\right>_{\mathbf{p}_{ij}},$ (41)
$\Rightarrow\tilde{D}_{j}^{\text{DA}}=\left<\gamma(\mathbf{R}_{ij}D_{j}^{\text{DA}}\mathbf{K}_{j}^{-1}\mathbf{p}_{j})_{z}+z_{ij}\right>_{\mathbf{p}_{ij}},$
where $\mathbf{R}_{ij},\mathbf{t}_{ij}]=\mathbf{E}_{j}^{-1}\mathbf{E}_{i}$, $z_{ij}=(\mathbf{t}_{ij})_{z}$, $\mathbf{p}_{ij}=\mathbf{R}_{ij}\mathbf{p}_{i}+\mathbf{t}_{ij}$, $\left$$\cdot\right$> indicates bilinear sampling.
We simplify above formulation as,
$\check{D}=\check{\gamma}D_{j}^{\text{DA}}+\check{z},$ (42)
where $\check{\gamma}$ and $\check{z}$ are ambiguous scale and shift, respectively. Similarly, we lift 2D pixels with this depth, yielding,
$\mathbf{P}=\check{D}\mathbf{K}^{-1}\mathbf{p}=\check{\gamma}D_{j}^{\text{DA}}\mathbf{K}^{-1}\mathbf{p}+\check{z}\mathbf{K}^{-1}\mathbf{p}.$ (43)
The offset vector can thus be calculated as follows.
$\mathbf{v}_{1}=\check{\gamma}D_{j,1}^{\text{DA}}\mathbf{K}^{-1}\mathbf{p}_{1}-\check{\gamma}D_{j}^{\text{DA}}\mathbf{K}^{-1}\mathbf{p}+\check{z}\mathbf{K}^{-1}(\mathbf{p}_{1}-\mathbf{p})$ (44)
Since $\mathbf{p}_{1}$ and $\mathbf{p}$ are neighborhood pixel and significantly close, the offset vector can further be approximated as,
$\mathbf{v}_{1}\approx\check{\gamma}(D_{j,1}^{\text{DA}}\mathbf{K}^{-1}\mathbf{p}_{1}-D_{j}^{\text{DA}}\mathbf{K}^{-1}\mathbf{p}),$ (45)
where $\mathbf{v}_{1}$ can again be expressed as a vector scaled by $\check{\gamma}$, analogous to the previous derivation with the estimated depth map in original target view.
In this manner, we eliminate the influence of scale and shift introduced by spatial dense depth reconstruction of DepthAnything, enabling the formulation of spatial surface normal vector map in target view and the computation of surface normal consistency loss $\tilde{\mathcal{L}}_{\text{SNC}}$.
Notably, as scale-ambiguous depth cannot be used for spatial warping to construct photometric losses, this spatial normal vector map derived from reconstructed dense depth can neither be obtained via forward warping, as pointing direction of elements are not coherent across view; nor via backward warping or modified forward–backward schemes, since camera extrinsics are calibrated in the metric physical world, whereas the depth output of DepthAnything remains scale-ambiguous.
VI-A4 Motion learning methods
In this work, we propose an adaptive joint motion learning strategy. Here, we present a detailed formulation of how prior works have realized surround-view motion learning.
FSM [4] estimates the pose of each camera independently. This motion learning paradigm can be written as,
$\{\boldsymbol{f}_{i}\}_{i=1}^{N}=\{\mathcal{P}_{\text{en}}(\mathbf{I}_{i}^{t},\mathbf{I}_{i}^{t^{\prime}})\}_{i=1}^{N},$ (46)
$\{\hat{\mathbf{T}}_{i}^{t\rightarrow t^{\prime}}\}_{i=1}^{N}=\{\mathcal{P}_{\text{de}}(\boldsymbol{f}_{i})\}_{i=1}^{N},$ (47)
where $\mathcal{P}_{\text{en}}$ and $\mathcal{P}_{\text{de}}$ indicates en- and decoder of pose network.
In addition, it enforces pose consistency by transforming all pose estimation to the coordinate of front camera as a global constraint. The pose transformation can be formulated as,
$\tilde{\mathbf{T}}_{i}^{t\rightarrow t^{\prime}}=\mathbf{E}_{1}^{-1}\mathbf{E}_{i}\hat{\mathbf{T}}_{i}^{t\rightarrow t^{\prime}}\mathbf{E}_{i}^{-1}\mathbf{E}_{1}$ (48)
where $\mathbf{E}_{1}$ indicates extrinsics of front camera. $\tilde{\mathbf{T}}_{i}^{t\rightarrow t^{\prime}}=[\tilde{\mathbf{R}}_{i}^{t\rightarrow t^{\prime}},\tilde{\mathbf{t}}_{i}^{t\rightarrow t^{\prime}}]$. Subsequently, it formulated pose consistency on translation and rotation separately as,
$\mathbf{t}_{\text{loss}}=\sum_{j=2}^{N}\|\hat{\mathbf{t}}_{1}^{t\rightarrow t^{\prime}}-\tilde{\mathbf{t}}_{1}^{t\rightarrow t^{\prime}}\|^{2},$ (49)
$\mathbf{R}_{\text{loss}}=\sum_{\varrho\in\{\phi,\theta,\psi\}}\sum_{j=2}^{N}\|\hat{\varrho}_{1}^{t\rightarrow t^{\prime}}-\tilde{\varrho}_{j}^{t\rightarrow t^{\prime}}\|^{2},$ (50)
$\mathcal{L}_{\text{PCC}}=\alpha_{t}\mathbf{t}_{\text{loss}}+\alpha_{r}\mathbf{R}_{\text{loss}},$ (51)
where $\alpha_{t}$ and $\alpha_{r}$ are weighting coefficients.
SurroundDepth [5] proposes a joint motion estimation strategy that aggregates feature maps from all cameras using a shared pose encoder and estimates a unified ego-motion in the LiDAR coordinate frame via a pose decoder. The pose of each individual camera is then recovered by distributing the joint motion through the calibrated extrinsic parameters.
$\hat{\mathbf{T}}_{i}^{t\rightarrow t^{\prime}}=\mathbf{E}_{i}^{-1}\mathcal{P}_{\text{de}}(\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{f}_{i})\mathbf{E}_{i},$ (52)
Both VFDepth [6] and CVCDepth [13] focus on estimating the front-camera motion. VFDepth conditions the pose decoder on aggregated multi-camera features, while CVCDepth relies solely on features from the front camera. The resulting motion estimation is formulated as,
$\hat{\mathbf{T}}_{i}^{t\rightarrow t^{\prime}}=\mathbf{E}_{i}^{-1}\mathbf{E}_{1}\mathcal{P}_{\text{de}}(\boldsymbol{f}_{1})\mathbf{E}_{1}^{-1}\mathbf{E}_{i},$ (53)
The alternative of motion learning adopted by CVCDepth [13] can be intuitively attributed to the observation that spatial structure within FoV of front camera are, in most driving scenarios, more sensitive to ego-motion and therefore provide informative cues for structure-from-motion (SfM) learning. However, this heuristic implicitly assumes a fixed dominance of the front view and overlooks the complementary motion cues available from other camera views. Therefore, in this work, we propose a adaptive joint motion learning strategy, in which the pose network adaptively learns and weights the contribution of each camera view for motion estimation.
VI-B Experiment Results
VI-B1 Implementation details
Evaluation metrics. The evaluation metrics used for our experiments are calculated as follows.
  • Absolute relative error (Abs Rel):
    $\frac{1}{n}\sum_{i\in n}|\hat{\mathbf{D}}(i)-\mathbf{D}(i)|/\mathbf{D}(i);$
  • Square relative difference (Sq Rel):
    $\frac{1}{n}\sum_{i\in n}\|\hat{\mathbf{D}}(i)^{2}-\mathbf{D}(i)^{2}\|/\mathbf{D}(i);$
  • Root mean square error (RMSE):
    $\sqrt{\frac{1}{n}\sum_{i\in n}\|\hat{\mathbf{D}}(i)-\mathbf{D}(i)\|^{2}};$
  • Root mean squared logarithmic error (RMSE log):
    $\sqrt{\frac{1}{n}\sum_{i\in n}\|\log\hat{\mathbf{D}}(i)-\log\mathbf{D}(i)\|^{2}}$
  • Accuracy with threshold ($\delta_{t}$):
    $\%~\text{of}~\hat{\mathbf{D}}(i)~\text{s.t.}~\max\left(\dfrac{\hat{\mathbf{D}}(i)}{\mathbf{D}(i)},\dfrac{\mathbf{D}(i)}{\hat{\mathbf{D}}(i)}\right)<1.25^{t},$
where $n$ indicates number of valid depths in groundtruth.
VI-B2 Addition experiment results
In Fig.7, we present additional examples of depth estimation, surface normal visualization, as well as pseudo-depth of DepthAnything and the corresponding surface normal maps computed from it on the DDAD and nuScenes datasets. These results further validate the effectiveness of our method in producing edge-aware, naturally transitioning, and smooth depth estimates under diverse conditions.
Figure 7: More comparison examples on the DDAD and nuScenes datasets are presented. Our method accurately estimates edge-aware, naturally transitioning, and smooth depth under diverse conditions.
In Fig.8, we present examples of point cloud reconstruction using estimated dense depth on the DDAD and nuScenes datasets, with comparisons against the baseline method CVCDepth [13]. The visualizations show that our proposed method produces geometrically regular point clouds with improved cross-view coherence and spatial consistency. For example, in Fig.8(a), lane markings are cleanly aligned across views, road lights stand upright above the ground plane, and distant vehicles are accurately projected and positioned. In contrast, in Fig.8(b), the baseline method incorrectly estimates the depth of a vehicle in the front view, resulting in erroneous 3D projection, while depth holes in the rear view further cause the vehicle to be wrongly projected onto the ground.
Figure 8: Examples of point cloud reconstruction comparison on DDAD and nuScenes datasets.

### v2

@@ III-C Geometric Feature Representation Enhancement
III-D Adaptive Joint Motion Learning
Pose estimation is a critical component for enabling pixel warping and subsequent view synthesis. Unlike prior approaches that estimate a joint motion in a fixed coordinate frame using features aggregated from all views, such as SurroundDepth and VFDepth, or that assume a fixed camera motion with view-specific features like CVCDepth, we propose an adaptive joint motion learning strategy. Our approach encourages the network to learn and emphasize informative cues for SfM learning, allowing pose estimation to adapt to varying view contributions.
Specifically, we leverage feature maps extracted from all cameras, denoted as $\boldsymbol{f}=\{\boldsymbol{f}_{i}\}_{i=1}^{N}$. We first apply spatial average pooling, followed by averaging across the camera dimension to obtain a global feature representation $\bar{\boldsymbol{f}}_{N}\in\mathbb{R}^{C}$ that aggregates holistic multi-view information. Instead of uniformly averaging camera features, we introduce a learnable fully-connected network $\xi$ to predict a weight vector $\boldsymbol{\omega}$, enabling adaptive emphasis on informative views for pose estimation. The architecture of the proposed motion learning module is illustrated in Fig.2(a) and is formulated as follows.
where $\mathcal{P}_{\text{de}}$ denotes pose decoder.
@@ IV-A Implementation Details
Training. Our networks were implemented in PyTorch and trained on four NVIDIA RTX 4090 GPUs. MonoViT-Small [24], modified and adapted to our surround-view setting, was employed as the depth network. ResNet-18 [9] was adopted as pose network following VFDepth [10]. During training, images from the previous and subsequent frames ($t^{\prime}\in{t-1,t+1}$) were used as temporal context. We trained the models using the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, a learning rate of $1\times 10^{-4}$, and 30/20 training epochs for DDAD/nuScenes dataset. A batch size of 1, consisting of images from 6 cameras, was used per GPU. Focal normalization [4] and the intensity alignment strategy proposed in VFDepth were applied during training (More details are provided in Section 2.1 of Supplementary material).
@@ IV-B Experiment Results
Monocular depth estimation. We first conduct a simple experiment on monocular depth estimation by altering the network architecture, modifying loss function of MonoViT to incorporate $\mathcal{L}_{\text{SNC}}$ and $\mathcal{L}_{\text{DSC}}$, and rename our method as GeoDepth. The qualitative results shown in Fig.5 demonstrate that our proposed method achieves more accurate, edge- and object-aware depth estimation. The quantitative results reported in Table I further show that GeoDepth outperforms MonoViT as well as the SoTA method EDS-Depth [22]. Notably, this performance is achieved efficiently and with minimal additional effort, by simply modifying the network architecture and incorporating additional loss terms, without introducing auxiliary modalities such as semantic segmentation, optical flows, or temporal cues. More detailed results are provided in Section 2.3 of the Supplementary material.
Figure 5: Visualization of depth estimation results on KITTI.
[…]
MonoViT 0.099 0.708 4.372 0.175 0.900 0.967 0.984
EDS-Depth 0.095 0.619 4.184 0.170 0.905 0.969 0.985
GeoDepth 0.094 0.641 4.175 0.167 0.906 0.970 0.986
TABLE I: Scale-ambiguous evaluation on KITTI eigen test split. Results best in bold, second best underlined).
Surround-view depth estimation. We compare our proposed method with other SoTA approaches. Quantitative evaluations for both metric and scale-ambiguous depth estimation on DDAD and nuScenes datasets are reported in Table II. Our method achieves substantially better or competitive performance and generalization capability compared with other baselines and DepthAnything V2 on both datasets, even when using ResNet-34 as depth network. Qualitative visualizations of surround-view depth estimation results on both datasets are presented in Fig.6. As shown, our method produces smooth, edge- and object-aware depth maps, significantly outperforming methods such as CVCDepth. Visualization of more examples as well as point cloud reconstruction on both datasets can be found in Section 2.4 and 2.5 of Supplementary material provided.
Figure 6: Visualization of depth estimation results on the DDAD and nuScenes datasets. White boxes indicate erroneous estimations.
@@
[…]
CVCDepth (Res34) [2] 0.270 5.273 15.127 0.400 0.603 0.810 0.895 0.289 3.503 7.469 0.353 0.642 0.848 0.923
GeoSurDepth (proposed) 0.193 3.049 13.227 0.307 0.719 0.885 0.944 0.246 2.688 6.885 0.318 0.694 0.877 0.937
TABLE II: Depth evaluation and generalization test results on DDAD and nuScenes datasets (* indicates reproduced results by VFDepth. Generalization tests were conducted using median-scaling due to different depth clipping ranges across two datasets.
@@ IV-C Ablation Studies
Adaptive joint motion learning. Table III shows improved depth estimation performance with proposed method against other baselines or methods for individual camera views and surround-view setting. We attribute this improvement to the learnable weighting of features extracted by the pose encoder, which enables the network to adaptively emphasize informative views and thereby achieve more effective SfM learning.
@@
TABLE III: Ablation study on different motion learning strategies on DDAD dataset (F, B, L, R implies front, back, left and right).
Spatial dense depth-based view synthesis. In this work, we leverage reconstructed spatial dense depth to further perform view synthesis and enforce surface normal consistency. As shown in Table IV, removing either $\tilde{\mathcal{L}}_{\text{SRC}}$ or $\tilde{\mathcal{L}}_{\text{SNC}}$ from the overall loss function leads to noticeable performance degradation. Fig.7 presents a qualitative example of view synthesis using both the estimated depth and the reconstructed spatial dense depth through spatial warping. The zoomed-in regions illustrate the complementary effect of these two depth sources on view synthesis and photometric penalization for 3D lifting.
@@
TABLE IV: Evaluation of the use of reconstructed spatial dense depth on metric depth estimation of DDAD dataset.
@@
Geometry guidance and enhancement via foundation models. Results reported in Table V demonstrate that enforcing 3D geometric accuracy through surface normal consistency, together with regularizing edge-aware depth estimation using 2D disparity gradients, leads to improved model performance. In contrast, removing CLIP from the depth network results in a slight drop in model performance. Fig.8 provides qualitative comparisons of these variants, illustrating less smooth depth transitions in low-texture regions, increased edge blurring at object boundaries when geometric guidance is absent, and reduced semantic details when geometric feature representation enhancement is removed.
@@
TABLE V: Ablation study on geometry guidance by DepthAnything V2 and feature representation enhancement by CLIP model.
@@ V Conclusion
In this work, we presented GeoSurDepth, a framework for self-supervised surround-view depth estimation that exploits geometry consistency as primary cue. By integrating vision foundation models as pseudo-geometry priors and for feature enhancement, enforcing 3D surface normal consistency, and regularizing object- and texture-level depth across views, GeoSurDepth achieves accurate and edge-aware depth estimations. A novel view synthesis pipeline provides additional photometric supervision through 2D-3D lifting and multi-contextual reconstruction, while an adaptive joint motion learning strategy enables the network to emphasize informative camera views for improved motion reasoning. Extensive experiments on KITTI, DDAD and nuScenes demonstrate that GeoSurDepth achieves SoTA performance, highlighting the importance of exploiting geometry coherence and consistency for robust self-supervised depth estimation.